OpenStudy (anonymous):

Evaluate -2(x + 1) for x = . Type a numerical answer

OpenStudy (redohawk):

-2(x+1) -2x - 2 x

OpenStudy (redohawk):

don't know, is this really all the info?

OpenStudy (anonymous):

yeah thats the whole question.

OpenStudy (anonymous):

After getting -2x-2=0 You can simplify to -2x=2 Then, -x=2/2 Next, -x=1 Finally, x=-1